A local restaurant serves German (or at least as close to it as you can get in Sydney) food prepared from local ingredients. They do, however, import several brands of beer (bier) directly from Germany. Currently the average price they pay for the beer is EUR37.56 per case. Last year the restaurant purchased 1,350 cases. The restaurant had revenues (net of other costs) of AUD147,800 and they expect that revenue to increase with the local rate of inflation, 4.1%, over the next year. The price of German beer is expected to increase at the rate of German inflation, 2.5%. The current exchange rate is AUD1.2513/EUR and remained unchanged from last year.

If the Euro experiences a real appreciation of 7.8% relative to the Australian dollar this year, and if the restaurant purchases the same amount of beer as last year, by how much will real profits change?

Select one:

a. -4.62%

b. -2.01%

c. -5.01%

d. -6.11%

e. -5.87%
Answer #1

Last year's Profit = Reveues - Cost of Beer bought = AUD 147,800 - 1,350 cases x EUR 37.56 per case x AUD 1.2513 / EUR = 147,800 - 1,350 x 37.56 x 1.2513 = AUD  84,352

This year's exhcnage rate = Last year's rate / (1 + real appreacition in EUR) = 1.2513 / (1 + 7.8%) = AUD  1.3489 /EUR

Hence, this year's profit (real profit, excluding impact of inflation) = Reveues - Cost of Beer bought = AUD 147,800 - 1,350 cases x EUR 37.56 per case x AUD 1.3489 / EUR = 147,800 - 1,350 x 37.56 x 1.3489 = AUD   79,403

Hence, cage in profit = Profit this year / profit last year - 1 =  79,403 / 84,352 -1 = -5.87%

Hence, the correct answer is the last option i.e option e. -5.87%